Shift report 20030222 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 22 2003 - 04:15:29 EST

  • Next message: Apache: "Shift report 20030222 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Feb 22 04:15:29 2003, Shift: 00:00-08:00
    Supervisor: MM, with: 
    Log book#14 pages 30 - 31
    
    Shift Summary:
    ------------------------------------------------------------
    1. Vacuum problem in yellow beampipe, meaning yellow beam can not be ramped properly. (It is lost before it is coged.)
    - Magnet powersupply problem in blue beam, causing frequent quench interlocks. Experts will come and try to fix this problem saturday morning
    2.Only pedestal runs taken
    3.No detectorproblems detected
    
    -Main control told me that it was certain that there would be no physics store before 8am this morning. As a consequence of this, I will leave the night shift and go to bed. (Flemming said this was ok, as long as magnets are turned to standby and TPC are turned down, wich HAS been done.)
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 22 2003 - 04:20:08 EST